# LLM Prompt Duel Optimizer: Efficient Label-Free Prompt Optimization

### v1

@@ 9      Expand $W,N$ with zero row/column for $p_{\text{new}}$;
Table 2: Test results of PDO when selecting the prompt with the highest development set accuracy (from the same candidate pools as in Table 1). The results are compared with supervised APO methods.
4 Experiment and Results
4.1 Experimental Setup
Datasets.
In this section, we conduct experiments to evaluate PDO on both closed-ended multiple-choice tasks and open-ended QA tasks. For the multiple-choice setting, we select 16 tasks from Big-Bench-Hard (Suzgun et al., 2022), using accuracy as the evaluation metric. For the open-ended QA setting, we consider four task categories from MS-MARCO (Bajaj et al., 2016), where the final evaluation metric is an integer score between 11 and 55, assigned by an LLM judge comparing the model’s output with the ground-truth answer provided by the original dataset.
LLM Judge Design.
For multiple-choice tasks, we require the LLM to produce both an answer and the accompanying reasoning given the instruction prompt. We then apply a dual-judge approach: if two prompts yield different answers, the LLM judge selects the prompt with the correct answer; if they yield the same answer, the judge decides based on the quality of reasoning. For open-ended tasks, the LLM judge design is more straightforward: each pair of responses is evaluated on accuracy, completeness, relevance, and clarity, consistent with the criteria of the final evaluation metric. To mitigate position bias, the order of the two outputs in each prompt pair is randomized before being fed into the judge template. A detailed rationale and analysis of the LLM preference judge design are provided in Section 5 and Appendix A.
(a) Description
(b) Entity
(c) Numeric
(d) Location
Baselines.
For the Big-Bench-Hard (BBH) dataset, we also evaluate PDO under a supervised setting to enable comparison with popular supervised APO methods, including APE (Zhou et al., 2022), OPRO (Yang et al., 2024), and Breeder (Fernando et al., 2023). The optimization itself remains label-free; labels are introduced only in the final stage to select the prompt that achieves the highest development-set accuracy among candidates generated through initial proposals and prompt mutations. Results for the without-label and with-label settings are reported in Tables 1 and 2, respectively.
4.2 Benchmark Results
Multiple-choice Tasks Performance.
As shown in Table 1, using only preference signals from the LLM judge and selecting the winner prompt by Copeland scores, PDO achieves the highest evaluation accuracy on 13/16 tasks. Notable gains include Tracking-7 (0.641 vs. 0.543, +9.8pp) and Web of Lies (0.942 vs. 0.861, +8.1pp). We examine the benefit of top-performer guided mutation in Appendix B.
In the labeled setting, when the final prompt is chosen by development-set accuracy, PDO remains competitive with state-of-the-art prompt optimization baselines, achieving the top score on 9/16 tasks. For the Geometric dataset, selecting the prompt based on development-set accuracy yields 0.712, outperforming all other methods, whereas selecting based on the LLM judge’s ranking results is only 0.598 accuracy, underperforming the label-free baseline SPO. In this particular case, this gap suggests that the LLM judge is less effective and noisier in identifying the best prompt on this dataset compared with others in BBH. We further examine the issue of LLM judge noise in Section 5.
Open-ended QA Tasks.
5 LLM Preference Judge Analysis
A fundamental bottleneck in PDO is that it depends on an LLM judge to approximate downstream performance through pairwise comparisons. In this section, we examine the factors that influence the effectiveness of the LLM judge in PDO.
Correlation Between Judge Noise and Task Performance.
We proxy LLM-judge effectiveness for the 16 BBH tasks by the test accuracy gap between Table 1 (prompt chosen by Copeland score, label-free) and Table 2 (prompt chosen by development-set accuracy). In our results, Tracking-7 and Web of Lies show small gaps and achieve top performance under both selection criteria, whereas Geometric exhibits a large discrepancy of 11.4 percentage points.
To investigate further, for each task we fix $|\mathcal{P}|=20$ instruction prompts with varying accuracy, run D-TS with the LLM judge, and compare it against an oracle judge that always selects the higher-accuracy prompt in each duel. At round $t$, we record the ground-truth accuracy rank of the current Copeland leader. Figure 3(a) shows that the oracle converges to the best prompt by round 4; with the LLM judge, Tracking-7 steadily improves to rank 2 and Web of Lies approaches rank $\approx 2.5$, while Geometric remains around ranks 6–8 across rounds. These trends confirm that judge reliability is closely related to the performance gaps observed in Tables 1 and 2.
Reducing Judge Noise.
We hypothesize that a key source of judge noise in BBH tasks is the inability to select the correct prompt when two prompts yield different answers. As an extension to the original label-free setting, we test whether providing a small fraction of labeled examples can help identify the top-ranked prompt.
(a) Judge noise across different BBH tasks.
(b) Noise reduction on Geometric with partial labels.
Figure 3: The accuracy of the LLM judge varies across tasks, with some tasks (e.g., Geometric) showing persistent misjudgments compared to an oracle judge. Introducing a fraction of true labels ($r\in\{0.3,0.5\}$) helps reduce the impact of noisy judgments and guide D-TS more quickly toward the top-ranked prompts.
Specifically, we assume that the ground-truth labels of an $r\in\{0.3,0.5\}$ fraction of examples are known in advance. On each round, if the chosen example $x_{t}$ belongs to this labeled subset, we determine the duel outcome using its true label: if prompt $i$ is correct and $j$ is incorrect, we update $W[i,j]\leftarrow W[i,j]+1$, and symmetrically for the reverse case. If both are correct or both are wrong, we add $0.5$ to each of $W[i,j]$ and $W[j,i]$. When $x_{t}$ is unlabeled, the LLM judge decides the duel as usual. Figure 3(b) shows that on Geometric, $r=0.3$ pulls the leader to roughly the top 3 by later rounds, while $r=0.5$ drives it close to ranks 1–2 with narrower ribbons. Overall, these results confirm our hypothesis: access to partial ground-truth labels mitigates systematic judge errors and steers D-TS toward the best prompt.
Impact of Judge Model Size.
We evaluate MS-MARCO performance when using Llama3.1-8B as the preference judge compared to the larger Llama3.3-70B (default). Table 3 reports the test accuracy of the prompt with the highest Copeland score at round 30. The comparison shows mixed but stable results: 70B slightly outperforms 8B on Entity and Numeric, while 8B surpasses 70B on Description and Location. Overall, both judges perform similarly and consistently outperform the baseline SPO.
Method Description Entity Numeric Location
SPO $4.385_{\pm\scriptscriptstyle 0.16}$ $4.242_{\pm\scriptscriptstyle 0.24}$ $3.981_{\pm\scriptscriptstyle 0.25}$ $4.292_{\pm\scriptscriptstyle 0.23}$
[…]
Table 3: Results on MS-MARCO using Llama-70B vs. 8B as the preference judge.
(a) Description
(b) Numeric
Figure 4: On the Numeric task, Llama-70B as judge enables D-TS to find strong prompts faster and reach higher scores, whereas on the Description task, Llama-8B slightly outperforms 70B overall.
Pairwise Preference vs. Pointwise Scoring.
6 Related Work
APO Methods: from Supervised to Label-free.
Traditional prompt optimization methods rely heavily on supervised signals from labeled validation sets, where the typical pipeline involves generating, mutating, and scoring prompts Zhou et al. (2022); Fernando et al. (2023); Pryzant et al. (2023); Guo et al. (2023); Schulhoff et al. (2024). More recently, studies have reduced supervision requirements Xiang et al. (2025); Chen et al. (2024); Madaan et al. (2023); Cheng et al. (2023); Dong et al. (2022). Notably, Self-Supervised Prompt Optimization (SPO) Xiang et al. (2025) eliminates external references by using output-vs-output comparisons, iteratively generating, executing, and comparing prompts through pairwise LLM judgments. However, SPO follows a greedy hill-climbing loop without a principled exploration–exploitation strategy, limiting its ability to allocate comparisons efficiently and to robustly identify high-performing prompts.
Bandit-Based Prompt Optimization.
7 Conclusion
In this paper, we introduce PDO, a prompt optimization method that reduces dependence on costly supervision by formulating label-free optimization as a dueling bandit problem guided by LLM preference feedback. By combining Double Thompson Sampling (D-TS) with top-performer–guided mutation, PDO adaptively searches for informative prompt comparisons to identify high-performing prompts. Experiments on multiple-choice BBH tasks and the open-ended MS-MARCO dataset show that PDO consistently achieves competitive performance against baseline methods across tasks. We further demonstrate that LLM judges provide reliable preference feedback overall, while also analyzing the impact of judge noise and the role of partial labels in reducing it during optimization.
Limitation
While our dueling-bandit framework provides a practical way to optimize prompts without ground-truth labels, it also has natural limitations. The quality of preference feedback depends on the meta prompt used to guide the LLM judge, and subtle differences in wording or criteria can influence its decisions. The judge’s own capability also matters: stronger foundation models tend to produce more reliable preferences, whereas smaller or less specialized models may introduce more noise, particularly in domain-specific tasks. Since the algorithm optimizes for the judge’s notion of quality rather than the true task metric, there is a risk of favoring stylistic patterns that the judge prefers. Although we include a thorough discussion and ablation studies of LLM judge design in Section 5 and Appendix A, due to computational resource constraints we were unable to conduct broader experiments across additional models. Despite these challenges, our formulation offers a principled and scalable starting point for label-free prompt optimization using preference feedback, and we hope it inspires future work on improving the alignment of LLM judge with task objectives.
@@ 7 Conclusion
Appendix A Further Discussion on LLM Judge Design
@@
A.1 Validating the Dual-Judge Approach
To further probe the dual-judge approach, we conduct the following experiments. For each dataset, we select two prompts with different accuracy levels: prompt AA as the higher-accuracy prompt and prompt $B$ as the lower-accuracy prompt. We then test the judge in two scenarios: (i) whether prompt A is preferred when A is correct and B is wrong, and (ii) whether prompt A is preferred when both A and B produce correct answers. The results are reported in Figure 6.
@@ A.1 Validating the Dual-Judge Approach
For most BBH tasks, we find strong evidence that the LLM judge reliably selects the higher-accuracy prompt when AA is correct and $B$ is wrong. We also observe moderate evidence that when both prompts produce the same answer, the judge prefers AA based on the quality of its reasoning. Datasets where this preference is consistent under both Answer and Reasoning conditions (Hyperbaton, Tracking-5, Tracking-7, and Web of Lies) show superior performance in Table 1. On the other hand, when the judge is inconsistent in identifying the higher-accuracy prompt AA in either condition, performance degrades (Geometric, Salient).
A.2 Weighted Preference Matrix Update
@@
Figure 6: For each BBH dataset, we construct a prompt pair $(A,B)$, where AA is the higher-accuracy prompt and $B$ is the lower-accuracy prompt. We report the proportion of comparisons in which the judge prefers AA. Answer (green) covers cases where AA is correct and $B$ is wrong; the judge evaluates by checking answers. Reasoning (blue) covers cases where both prompts produce the correct answer; the judge compares their reasoning chains. Aggregate pools all examples across datasets.
@@
Takeaway.
A.3 Case Study
Appendix B The Effect of Prompt Mutation
PDO extends the standard dueling-bandit framework with top-performer–guided mutation, which incrementally expands the initial candidate pool by mutating top-performing prompts. We demonstrate its effectiveness on two BBH datasets: Web of Lies and Tracking-7. Starting from $|\mathcal{P}|=20$ instructions, D-TS runs pairwise duels. For the mutate setting, at rounds 10 and 20 we prune the 10 lowest–Copeland-score prompts and generate 10 new candidates by mutating the current Copeland leader, then continue dueling with the updated pool. In contrast, the non_mutate baseline keeps the initial 20 prompts fixed and selects leaders only from this pool. At each round we report the ground-truth accuracy of the current Copeland leader. As shown in Figure 8, the two curves coincide before round 10, after which mutation consistently improves accuracy on both Web of Lies and Tracking-7, moving beyond the limits of the original pool and steering the search toward stronger regions of the prompt space.
Appendix C Experiment Details
C.1 Hyperparameters
For the main experiments in Section 4.2, we initialize with 20 candidate prompts for BBH and 50 candidate prompts for MS-MARCO. Each experiment runs for 30 rounds, with 50 duels per round. At rounds 10 and 20, we apply prompt mutation by selecting the top-3 prompts ranked by Copeland scores and generating 10 new prompts. At the same time, we prune the 10 lowest-ranked prompts by Copeland scores. We always select the prompt with the highest Copeland score as the winner, using the average win rate as a tiebreaker when multiple prompts share the same score. The D-TS parameter is fixed at $\alpha=1.2$ throughout all experiments.
C.2 Prompt generation strategies
C.3 Baselines
SPO follows an iterative optimize–execute–evaluate loop, where new prompts are generated, executed, and then compared through pairwise LLM judgments of their outputs. For supervised APO methods, the typical pipeline involves generating, mutating, and scoring prompts. APE frames prompt design as black-box optimization, generating candidate instructions from input–output demonstrations and refining them through iterative search. OPRO treats the LLM itself as the optimizer: a meta-prompt with task descriptions and prior results guides the model to propose and evaluate new candidates in an iterative loop. Breeder applies an evolutionary approach, jointly evolving task-prompts and mutation-prompts through LLM-driven mutation and selection.
C.4 Datasets
BIG-bench Hard (BBH).
MS MARCO.
Appendix D Computational Complexity in # of LLM Calls
We analyze complexity purely in terms of the number of LLM calls, assuming each call has identical unit cost. All previously generated prompts’ predictions are cached and reused across rounds.
Notation.
$P$: total number of prompts produced (including the initial baselines); $P_{0}$: number of initial baselines (constant); $R$: number of optimization rounds (one new prompt per round, so $P=P_{0}+R$); $B$: number of examples used to score a new prompt per round (for caching its predictions); $M$: number of duels executed per round; NN: number of labeled examples used to score each prompt in supervised APO. We treat mutation cost as $\mathcal{O}(1)$ and omit it.
Supervised APO with labels.
Each generated prompt is scored once on NN labeled examples:
$\text{Calls}_{\text{sup}}\;=\;\mathcal{O}(P\,N).$
PDO (duels + caching).
At each round, we (i) predict for the new prompt on $B$ examples, incurring $B$ calls, and (ii) run $M$ duels using cached predictions for older prompts, incurring $M$ calls. Thus, the per-round cost is $B+M$ calls. Over $R$ rounds (equivalently, $P-P_{0}$ new prompts),
$\text{Calls}_{\text{PDO}}\;=\;\mathcal{O}\!\big(R\,(B+M)\big)\;=\;\mathcal{O}\!\big(P\,(B+M)\big).$
Comparison.
[…]
Remarks.
If $B\ll N$, PDO can require fewer calls than supervised APO because each new prompt is only partially evaluated and preferences guide the search. If $B=N$ (new prompts fully evaluated as in supervised APO), then
$\text{Calls}_{\text{PDO}}\;=\;\mathcal{O}(P\,(N+M)),$
which is less efficient than supervised APO due to the additional $M$ duels per round. Hence, PDO is advantageous in label-scarce regimes ($B<N$), while supervised APO is preferable in label-rich regimes ($B=N$).

### v2

@@ Appendix D Discussion on Evaluation Scope
Table 6: We report test results for PDO when selecting the prompt with the highest development-set accuracy from the same candidate pools as in Table 1. This setting serves as a proxy for replacing the LLM preference judge with an oracle. We compare these results against supervised APO methods and find that PDO remains competitive with state-of-the-art prompt optimization baselines, achieving the best performance on 9 out of 16 tasks.
(a) Causal
(b) Tracking-7
(c) Web of Lies
Figure 7: Test performance when using Claude 4.5 vs. Llama 3.3 as the LLM preference judge on BBH tasks under the same judge-call budget. Using Claude 4.5 more reliably identifies higher-performing prompts; the improvement is largest on tasks where the Llama 3.3 judge is noisier (e.g., Causal in Figure 7a and Geometric in Figure 3(b)).
Dataset Non-mutate Mutate Diff. $p$-value
Causal 0.709 0.703 -0.006 >0.05>0.05
Date 0.927 0.938 -0.011 0.0291
DisambigQA 0.766 0.760 -0.006 >0.05>0.05
Formal 0.759 0.767 -0.008 >0.05>0.05
Geometric 0.611 0.672 -0.061 $9.7\times 10^{-5}$
Hyperbaton 0.935 0.943 -0.008 >0.05>0.05
Logical-5 0.821 0.831 -0.010 >0.05>0.05
Logical-7 0.755 0.750 -0.005 >0.05>0.05
Navigate 0.903 0.917 -0.014 0.0146
Penguins 0.943 0.955 -0.012 0.0236
Salient 0.719 0.711 -0.008 >0.05>0.05
Snarks 0.863 0.876 -0.013 >0.05>0.05
Tracking-5 0.793 0.833 -0.040 0.0252
Tracking-7 0.585 0.647 -0.062 0.0040
Tracking-3 0.949 0.957 -0.008 0.0491
Web of Lies 0.919 0.979 -0.060 $1.5\times 10^{-6}$
Table 7: Effect of prompt mutation on BBH. We compare the final Copeland-winner prompt selected by PDO when mutation is disabled (Non-mutate) versus when mutation is enabled (Mutate). For each BBH task, we report the mean test accuracy over 10 independent runs. We perform a two-sample $t$-test per task to assess whether mutation yields a statistically significant accuracy change. Overall, 8 out of 16 tasks show significant improvement at $p<0.05$. We visualize representative tasks (Web of Lies and Tracking-7) in Figure 4. Tasks with small negative differences show only minor changes and are likely attributable to run-to-run noise; across all tasks considered, mutation does not materially degrade performance.
@@
Appendix E Further Discussion on LLM Judge Design for Closed-Ended Tasks
@@
E.1 Validating the Dual-Judge Approach
To further probe the dual-judge approach, we conduct the following experiments. For each dataset, we select two prompts with different accuracy levels: prompt $A$ as the higher-accuracy prompt and prompt $B$ as the lower-accuracy prompt. We then test the judge in two scenarios: (i) whether prompt A is preferred when A is correct and B is wrong, and (ii) whether prompt A is preferred when both A and B produce correct answers. The results are reported in Figure 8.
@@ E.1 Validating the Dual-Judge Approach
For most BBH tasks, we find strong evidence that the LLM judge reliably selects the higher-accuracy prompt when $A$ is correct and $B$ is wrong. We also observe moderate evidence that when both prompts produce the same answer, the judge prefers $A$ based on the quality of its reasoning. Datasets where this preference is consistent under both Answer and Reasoning conditions (Hyperbaton, Tracking-5, Tracking-7, and Web of Lies) show superior performance in Table 1. On the other hand, when the judge is inconsistent in identifying the higher-accuracy prompt $A$ in either condition, performance degrades (Geometric, Salient).
E.2 Weighted Preference Matrix Update
@@
Takeaway.
E.3 Case Study
Figure 8: For each BBH dataset, we construct a prompt pair $(A,B)$, where $A$ is the higher-accuracy prompt and $B$ is the lower-accuracy prompt. We report the proportion of comparisons in which the judge prefers $A$. Answer (green) covers cases where $A$ is correct and $B$ is wrong; the judge evaluates by checking answers. Reasoning (blue) covers cases where both prompts produce the correct answer; the judge compares their reasoning chains. Aggregate pools all examples across datasets.